# Cobordism of Domes Over Curves

### v1

@@ 4. Not Every Integral Curve is Cobordant to a Unit Rhombus
This section proves that not every integral curve is orientably cobordant to a unit rhombus. Our method is to generalize [1, Theorem 4.3] to allow for graph surfaces of arbitrary genus. We rephrase a cobordism as a sample polyhedron with multiple boundary components, and project down to smaller moduli space to account for any possible polygon to be chosen for the cobordism.
@@
Note that the skew symmetric form on $\mathbb{E}^{\textbf{P}}/\text{Isom}^{+}\mathbb{E}$ is induced by the skew symmetric form $\omega$ on $\mathbb{E}^{\textbf{P}}/\mathbb{E}$ after taking the quotient by the action of $SO(3)^{\times n}$. We show that the pullback of this skew symmetric form $\omega^{\prime}$ is null. Also note that the orientability hypothesis is necessary here, see §$\S$ 5.3
@@
The proof of [1, Theorem 4.3] follows word for word to show that $\delta:\mathbb{E}^{S}/\mathbb{E}\to\mathbb{E}^{\textbf{P}}/\mathbb{E}$ is isotropic by a series of collapses. To show that this survives to quotient by $SO(3)^{\times n}$, we use Lemma 3.4. Because $T_{\textbf{p}}SO(3)^{\times n}\textbf{p}=\ker\omega_{\textbf{p}}$, taking the quotient by $SO(3)^{\times n}$ on $\mathbb{E}^{\textbf{P}}/\mathbb{E}$ gives a new nondegenerate skew-symmetric form $\omega^{\prime}$ on $\mathbb{E}^{\textbf{P}}/\text{Isom}^{+}\mathbb{E}$, and the pullback of the form to $\mathbb{E}^{S}/\text{Isom}^{+}\mathbb{E}$ is still null. ∎
@@
Originally, Glazyrin and Pak only considered integral curves with one connected component. So Theorem 1.2 is not an answer to Conjecture 1.1 in the orientable case, but we give one in the following corollary.
Corollary 4.3.
There exists an integral curve $\gamma$ with one connected component that is not orientably coboardant to any unit rhombus $\rho$.
@@
Choose a pair of unit rhombi $\rho,\rho^{\prime}\in\mathcal{M}_{4}$ such that $\rho\cup\rho^{\prime}$ is not cobordant to any third unit rhombus $\rho^{\prime\prime}\in\mathcal{M}_{4}$. Let $\rho=[v_{1}v_{2}v_{3}v_{4}]$ and $\rho^{\prime}=[v_{1}^{\prime}v_{2}^{\prime}v_{3}^{\prime}v_{4}^{\prime}]$. Position $\rho$ and $\rho^{\prime}$ so that $v_{1}=v_{1}^{\prime}$, and the distances $|v_{2},v_{2}^{\prime}|=|v_{4},v_{4}^{\prime}|=1$. By adding the unit triangles $[v_{1}v_{2}v_{2}^{\prime}]$ and $[v_{1}v_{4}v_{4}^{\prime}]$, we see that $\rho\cup\rho^{\prime}$ is coboardant to the perimeter $\gamma=[v_{4}v_{3}v_{2}v_{2}^{\prime}v_{3}^{\prime}v_{4}^{\prime}]$. Therefore, our choice of $\rho,\rho^{\prime}$ implies that $\gamma$ is not cobordant to any unit rhombus, and $\gamma$ has only one connected component. ∎
Glazyrin and Pak conjectured that a positive answer to Conjecture 1.1 would involve showing a cobordism between an integral curve and finitely many unit rhombi, and then a cobordism between two unit rhombi and one unit rhombus, see [4, Conj. 5.15, Conj. 5.16]. Thus Theorem 1.2 is a negative answer to [4, Conj. 5.16] and Corollary 4.3 is a negative answer to Conjecture 1.1, both in the orientable case.
@@ 5. Final remarks
The proof of Theorem 1.2 does not account for the fact that the unit rhombi in question can be oriented in some manner with respect to each other. That is, we have chosen to define $\mathbb{E}^{\textbf{P}}/\text{Isom}^{+}\mathbb{E}$ with the group of orientation preserving isometries $\text{Isom}^{+}\mathbb{E}$ acting on each polygon separately. However, we could consider $\text{Isom}^{+}\mathbb{E}$ as acting on all of the spaces together, so instead of $\mathbb{E}^{P_{1}}/\text{Isom}^{+}\mathbb{E}\times\dotsc\times\mathbb{E}^{P_{n}}/\text{Isom}^{+}\mathbb{E}$ we would consider $\mathbb{E}^{P_{1}}/\text{Isom}^{+}\mathbb{E}\times\mathbb{E}^{P_{2}}\times\dotsc\times\mathbb{E}^{P_{n}}$. This approach could answer more general questions. For example, Glazyrin and Pak conjectured that there are two unit triangles $\Delta_{1},\Delta_{2}\subset\mathbb{E}$ which are not cobordant, see [4, Conjecture 5.13]. In this question, the relative translation and rotation of the triangles is important, so our current techniques are insufficient.
Additionally, consider Steinhaus’ 1957 problem on tetrahedral chains, see [7]. A tetrahedral chain is a polyhedra constructed by attaching regular tetrahedra along faces to form a chain. These can be viewed as cobordisms between two triangles that satisfy stricter conditions than general domes. Steinhaus asked if tetrahedral chains can be closed, and if they are dense in $\mathbb{E}$. In contrast to general domes, the first question was shown to be false by Swierczkowski in [9]. However, the second question remains open. Recently, Stewart showed in [8] that the group generated by reflections across the faces of a regular tetrahedra is dense in $SO(3)$, but this does not resolve Steinhaus’ question because it does not consider the translations in the full group $SO(3)\rtimes\mathbb{E}$. I.e., Stewart’s approach considers only the relative rotation, not the relative translation of the two triangles.
@@
The minimal number $k$ of unit rhombi needed for the cobordism in Theorem 1.3 gives a measure of complexity for an integral curve $\gamma$. In contrast to Conjecture 1.1 which proposed that $k=1$ for all $\gamma$, we conjecture that $k=\Theta(|\gamma|)$. Theorem 1.3 gives a quadratic upper bound $k=O(|\gamma|^{2})$. We do not believe this is optimal. Pak proposes a modified proof of Theorem 1.3 which uses a reduction to generic integral curves rather than planar integral curves to improve to a linear bound. [^1] I. Pak, Personal communication. (See [4, §2] for terminology.) We also conjecture that the construction in Corollary 4.3 can be extended to arbitrarily many rhombi to prove that $k=\Omega(|\gamma|)$.
@@
The orientability hypothesis in Theorem 4.1 is necessary for the cancellation argument, see [1,  Remark 2.3] for more detail, and [1, Remark 4.4] for a counterexample with a non-orientable surface. This limits the techniques used in this paper to the orientable case, and new techniques may be necessary to study non-orientable cobordisms.
@@ Acknowledgements
I am grateful to Alexey Glazyrin and Dmitrii Korshunov, for helpful comments and clarifications. I am also thankful to Igor Pak for introducing me to this problem and for careful reading of my drafts.

### v2

@@ 4. Not Every Integral Curve is Cobordant to a Unit Rhombus
This section proves that not every integral curve is orientably cobordant to a unit rhombus. Our method is to generalize [AK24, Theorem 4.3] to allow for graph surfaces of arbitrary genus. We rephrase a cobordism of a certain combinatorial type as a sample polyhedron with multiple boundary components, and project down to the smaller moduli space with one fewer boundary component to account for any possible unit rhombus to be chosen for the cobordism.
@@
Note that the skew symmetric form on $\mathbb{E}^{\textbf{P}}/\text{Isom}^{+}\mathbb{E}$ is induced by the skew symmetric form $\omega$ on $\mathbb{E}^{\textbf{P}}/\mathcal{E}$ after taking the quotient by the action of $SO(3)^{\times n}$. We show that the pullback of this skew symmetric form $\omega^{\prime}$ is null. Also note that the orientability hypothesis is necessary here, see §$\S$ 5.3
@@
The proof of [AK24, Theorem 4.3] follows word for word to show that $\delta:\mathbb{E}^{S}/\mathcal{E}\to\mathbb{E}^{\textbf{P}}/\mathcal{E}$ is isotropic by a series of collapses. To show that this survives in the quotient by $SO(3)^{\times n}$, we use Lemma 3.4. Because $T_{\textbf{p}}SO(3)^{\times n}\textbf{p}=\ker\omega_{\textbf{p}}$, taking the quotient by $SO(3)^{\times n}$ on $\mathbb{E}^{\textbf{P}}/\mathcal{E}$ gives a new nondegenerate skew-symmetric form $\omega^{\prime}$ on $\mathbb{E}^{\textbf{P}}/\text{Isom}^{+}\mathbb{E}$, and the pullback of the form to $\mathbb{E}^{S}/\text{Isom}^{+}\mathbb{E}$ is still null. ∎
@@
Glazyrin and Pak did not consider unions of integral curves, so Theorem 4.3 is not an answer to Conjecture 1.1 in the orientable case, but we give one in the following.
Theorem 1.2.
There exists an integral curve $\gamma$ of length $5$ that is not orientably cobordant to any unit rhombus $\rho$.
@@
Choose a pair of unit rhombi $\rho,\rho^{\prime}\in\mathcal{M}_{4}$ such that $\rho\cup\rho^{\prime}$ is not cobordant to any third unit rhombus $\rho^{\prime\prime}\in\mathcal{M}_{4}$. Let $\rho=[v_{1}v_{2}v_{3}v_{4}]$ and $\rho^{\prime}=[v_{1}^{\prime}v_{2}^{\prime}v_{3}^{\prime}v_{4}^{\prime}]$. Position $\rho$ and $\rho^{\prime}$ so that $v_{1}=v_{1}^{\prime}$, $v_{2}=v_{2}^{\prime}$, and the distance $|v_{3},v_{3}^{\prime}|=1$. Note that from our definition of $\mathbb{E}^{\rho,\rho^{\prime}}/\text{Isom}^{+}\mathbb{E}$, it does not matter how $\rho$ and $\rho^{\prime}$ are oriented relative to each other, see §5.1. By adding the unit triangle $[v_{2},v_{3},v_{3}^{\prime}]$, we see tht $\rho\cup\rho^{\prime}$ is cobordant to the perimeter $\gamma=[v_{1}v_{4}v_{3}v_{3}^{\prime}v_{4}^{\prime}]$. Therefore, our choice of $\rho,\rho^{\prime}$ implies that $\gamma$ is not cobordant to any unit rhombus, and $\gamma$ has only one connected component. ∎
Glazyrin and Pak conjectured that a positive answer to Conjecture 1.1 would involve showing a cobordism between an integral curve and finitely many unit rhombi, and then a cobordism between two unit rhombi and one unit rhombus, see [GP22, Conj. 5.15, Conj. 5.16]. Thus Theorem 4.3 is a negative answer to [GP22, Conj. 5.16] and Theorem 1.2 is a negative answer to Conjecture 1.1, both in the orientable case.
@@ 5. Final remarks
The proof of Theorem 4.3 does not account for the fact that the unit rhombi in question can be oriented in some manner with respect to each other. That is, we have chosen to define $\mathbb{E}^{\textbf{P}}/\text{Isom}^{+}\mathbb{E}$ with the group of orientation preserving isometries $\text{Isom}^{+}\mathbb{E}$ acting on each polygon separately. However, we could consider $\text{Isom}^{+}\mathbb{E}$ as acting on all of the spaces together, so instead of $\mathbb{E}^{P_{1}}/\text{Isom}^{+}\mathbb{E}\times\dotsc\times\mathbb{E}^{P_{n}}/\text{Isom}^{+}\mathbb{E}$ we would consider $\mathbb{E}^{P_{1}}/\text{Isom}^{+}\mathbb{E}\times\mathbb{E}^{P_{2}}\times\dotsc\times\mathbb{E}^{P_{n}}$. This approach could answer more general questions. For example, Glazyrin and Pak conjectured that there are two unit triangles $\Delta_{1},\Delta_{2}\subset\mathbb{E}$ which are not cobordant, see [GP22, Conjecture 5.13]. In this question, the relative translation and rotation of the triangles is important, so our current techniques are insufficient.
Additionally, consider Steinhaus’ 1957 problem on tetrahedral chains, see [Stei57]. A tetrahedral chain is a polyhedron constructed by attaching regular tetrahedra along faces to form a chain. These can be viewed as cobordisms between two triangles that satisfy stricter conditions than general domes. Steinhaus asked if tetrahedral chains can be closed, and if they are dense in $\mathbb{E}$. In contrast to general domes, the first question was shown to be false by Swierczkowski in [Świ59]. However, the second question remains open. Recently, Stewart showed in [Stew19] that the group generated by reflections across the faces of a regular tetrahedra is dense in $SO(3)$, but this does not resolve Steinhaus’ question because it does not consider the translations in the full group $SO(3)\rtimes\mathbb{E}$. I.e., Stewart’s approach considers only the relative rotation, not the relative translation of the two triangles.
@@
The minimal number $k$ of unit rhombi needed for the cobordism in Theorem 1.3 gives a measure of complexity for an integral curve $\gamma$. In contrast to Conjecture 1.1 which proposed that $k=1$ for all $\gamma$, we conjecture that $k=\Theta(|\gamma|)$. Theorem 1.3 gives a quadratic upper bound $k=O(|\gamma|^{2})$. We do not believe this is optimal. For instance, the proofs of Lemma 2.1, Lemma 2.2 do not use any triangles in the rhombus equivalences; only rhombi are added. And the proofs could be optimized to improve the coefficients of the bound, for instance, by choosing three points in the plane $H$ rather than two in the proof of Lemma 2.1, but this would still give a quadratic bound in $|\gamma|$. Additionally, Pak proposes a modified proof of Theorem 1.3 which uses a reduction to generic integral curves rather than planar integral curves to improve to a linear bound. [^1] I. Pak, Personal communication. (See [GP22, §2] for terminology.) We also conjecture that the construction in Corollary 1.2 can be extended to arbitrarily many rhombi to prove that $k=\Omega(|\gamma|)$.
@@
The orientability hypothesis in Theorem 4.1 is necessary for the cancellation argument, see [AK24,  Remark 2.3] for more detail, and [AK24, Remark 4.4] for a counterexample with a non-orientable surface. This limits the techniques used in this paper to the orientable case, and new techniques may be necessary to study non-orientable cobordisms.
@@ Acknowledgements
I am grateful to Alexey Glazyrin and Dmitrii Korshunov, for helpful comments and clarifications. I am also thankful to Igor Pak for introducing me to this problem and for careful reading of my drafts. Finally, I would like to thank multiple anonymous referees for insightful comments and suggestions.